# MERaLiON-SER: Robust Speech Emotion Recognition Model for English and SEA Languages

### v1

@@ 1 Introduction
Human speech is a rich medium not only for semantic communication, but also for conveying emotional and paralinguistic information — such as affective states, intention, and interpersonal stance. Recognising emotions from speech has become a foundational task in affective computing and human-machine interaction: accurate speech emotion recognition (SER) enables applications ranging from virtual assistants that adapt to user mood, to mental-health monitoring, to empathetic conversational agents in service and robotics domains.
@@ 9 Conclusions
We extend our sincere gratitude to A*STAR management to support this work. The computational work for this article was fully performed on resources of the National Supercomputing Centre (NSCC), Singapore (https://www.nscc.sg). We also thank student intern Arjun Srinivas to help for data and automated emotion labeling work. This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore.

### v2

@@ 1 Introduction
Human speech is a rich medium not only for semantic communication, but also for conveying emotional and paralinguistic information such as affective states, intention, and interpersonal stance. Recognising emotions from speech has become a foundational task in affective computing and human-machine interaction: accurate speech emotion recognition (SER) enables applications ranging from virtual assistants that adapt to user mood, to mental-health monitoring, to empathetic conversational agents in service and robotics domains.
@@ 9 Conclusions
The computational work for this article was fully performed on resources of the National Supercomputing Centre (NSCC), Singapore (https://www.nscc.sg). We acknowledge Frank and his team from Asiastar International Consultancy Pte Ltd, who were engaged as an external consultancy to perform human annotations for the speech emotion annotation task. We also thank student intern Arjun Srinivas to help for data and automated emotion labeling work. This research is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.